solve b^4 + 2b^2 -24 =0
thank you

Answer by scott8148(6628)

this is a quadratic in b^2 ___ (b^2)^2 + 2(b^2) - 24 = 0

factoring ___ (b^2 + 6)(b^2 - 4) = 0 ___ (b^2 + 6)(b + 2)(b - 2) = 0

b = ±2, ±i√6